# Strategies for Pretraining Neural Operators

### v1

@@ Abstract
Pretraining for partial differential equation (PDE) modeling has recently shown promise in scaling neural operators across datasets to improve generalizability and performance. Despite these advances, our understanding of how pretraining affects neural operators is still limited; studies generally propose tailored architectures and datasets that make it challenging to compare or examine different pretraining frameworks. To address this, we compare various pretraining methods without optimizing architecture choices to characterize pretraining dynamics on different models and datasets as well as to understand its scaling and generalization behavior. We find that pretraining is highly dependent on model and dataset choices, but in general transfer learning or physics-based pretraining strategies work best. In addition, pretraining performance can be further improved by using data augmentations. Lastly, pretraining is additionally beneficial when fine-tuning in scarce data regimes or when generalizing to downstream data similar to the pretraining distribution. Through providing insights into pretraining neural operators for physics prediction, we hope to motivate future work in developing and evaluating pretraining methods for PDEs.
@@ 1 Introduction
11footnotetext: These authors contributed equally to this work**footnotetext: Corresponding Author
@@
Neural operators for PDEs have gained substantial interest in recent years due to their ability to quickly predict physics through inference (Li et al., 2021; Lu et al., 2021a; Brandstetter et al., 2023). Despite potential speed gains, neural operators currently struggle to generalize to unseen physics, and initial training can be slow (Lu et al., 2022; Gupta & Brandstetter, 2022). To address this issue, many works have explored different strategies to improve generalization by incorporating additional system information (Lorsung et al., 2024; Liu et al., 2023; Takamoto et al., 2023) and pretraining neural operators across large, diverse physics to quickly fine-tune to solve PDEs (Hao et al., 2024; McCabe et al., 2023; Shen et al., 2024; Hang et al., 2024; Goswami et al., 2022). Despite showing good performance, these works usually require the use of tailored neural operators and datasets to learn different physics. This contrasts with broader deep learning trends in which pretraining methods can universally benefit models; for example, pretraining losses that are applied across CNN models (Noroozi & Favaro, 2016; Lee et al., 2017) or GNN models (Hu et al., 2020). As a result, in this work, we consider existing pretraining frameworks, as well as propose novel methods for pretraining PDE models that are flexible and can be applied across architectures or datasets.
@@
Through this work, we hope to broaden the understanding of how neural operators can be pretrained for physics prediction. We organize existing pretraining strategies, propose novel vision-inspired strategies, and include common pretraining baselines to assemble a broad set of methods for learning PDE representations. We find that PDE pretraining varies depending on model and dataset choice, but in general using transfer or physics-based pretraining strategies work well. In addition, transformer or CNN-based architectures tend to benefit more from pretraining than vanilla neural operators. Furthermore, the use of data augmentations consistently improves pretraining performance in different models, datasets, and pretraining strategies. Lastly, we find that pretraining is more beneficial when fine-tuning in low-data regimes, or when downstream data is more similar to pretraining data. We hope that these insights can be used to guide future work in the development and evaluation of pretraining methods for PDEs. We open source our code here: https://github.com/anthonyzhou-1/pretraining_pdes .
@@ 2 Related Works
The field of neural operators has grown rapidly in recent years, with many architectures developed to accurately approximate solutions to PDEs (Li et al., 2021; 2023a; Gupta & Brandstetter, 2022; Brandstetter et al., 2023; Lu et al., 2021a). Many works expanded on this to propose architectures to solve PDEs more quickly, with less compute, or on irregular grids (Li et al., 2023b; Hemmasian & Barati Farimani, 2023; Li et al., 2023c), and as a result, within a range of test problems, neural operators can solve PDEs quickly and accurately. However, neural operators still struggle to generalize across diverse physics, and as a result many approaches have been developed to pretrain neural operators. We summarize these past works in Table 1, and briefly describe the main approaches here.
@@ 2.1 PDE Transfer Learning
Category PDEs Characteristic Reference
Transfer Darcy, Elasticity Fine-tuning task layers to transfer between domains/dynamics Goswami et al. (2022)
Poisson, INS, Wave, FP Design of a transferrable model through reparameterizing neurons Zhang et al. (2023c)
Heat, Adv, Nag, Burg, NS, AC Combining operator modules with gating for different PDEs Tripura & Chakraborty (2023)
NS, Elasticity Using variable pos. encoding and masked pretraining across domains Rahman et al. (2024)
Large Models Poisson, Helm Scaling model and dataset size to characterize transfer behavior Subramanian et al. (2023)
SWE, DiffReact, CNS Embedding PDEs to a common space and using an Axial ViT McCabe et al. (2023)
INS, CNS, SWE, DiffReact Using denoising and Fourier attention with large models and datasets Hao et al. (2024)
INS, CNS, SWE, DiffReact Training a conditional transformer across large PDE datasets Hang et al. (2024)
Poisson, Helm, NS, Wave, AC Scaling operator transformers to large, diverse datasets Herde et al. (2024)
Contrastive KdV, Burg, KS, INS Using Lie Symmetries to in self-supervised contrasive learning Mialon et al. (2023)
Heat, Advection, Burg Using physics-informed distance metrics in a contrastive framework Lorsung & Farimani (2024)
Burg, Adv-Diff, NS Using physical invariances to contrastively learn an encoder Zhang et al. (2023b)
Meta-Learning HGO, Elasticity, Tissue Using a model-agnostic meta-learning loss to learn across tasks Zhang et al. (2023a)
LV, GS, NS Using a novel loss term to maximize shared learning between PDEs Yin et al. (2021)
LV, GS, GO, NS Using a hyper-network to adapt PDE operators for specific tasks Kirchmeyer et al. (2022)
In-Context Poisson, Helm, DiffReact, NS Evaluating masked pretraining and in-context learning for PDEs Chen et al. (2024)
Poisson, DiffReact In-context learning for PDEs through prompting a transformer Yang et al. (2023)
@@ 2.2 Large PDE Modeling
An extension of transfer learning is to train large models on diverse physics datasets, with the intention of learning transferable representations through scaling behavior (Wei et al., 2022; Kaplan et al., 2020; Brown et al., 2020). 54 initially explores this scaling behavior by training large neural operator models on large PDE datasets to evaluate its ability to adapt to different coefficients. McCabe et al. (2023) propose a tailored architecture for solving problems across different physics, and Hao et al. (2024) expand on this by making architectural advancements and training on more diverse physics. Despite different approaches and datasets, these works generally rely on tailored, scalable architectures for large PDE datasets; pretraining is framed as physics prediction across diverse physics and fine-tuning is done on the pretraining distribution or on unseen coefficients/PDEs.
@@ 3.1.2 Physics-Agnostic Data Augmentations
In computer vision literature, many successful data augmentations heavily modify inputs (Chen et al., 2020); in particular, cropping and cutting out portions of an image would not respect physics if adapted to the PDE domain. Following this, we investigate the effect of data augmentations that are physics-agnostic, in that they can be applied to any PDE since the augmentation does not preserve the underlying dynamics. Following recent work on denoising neural operator architectures (Hao et al., 2024), we consider adding Gaussian noise during pretraining (Noise). Furthermore, we consider scaling the PDE solution (Scale), an approach similar to a color distortion, in which the PDE solution values are multiplied by a random constant. For certain simple PDEs, scaling can preserve physics, but this is not generally true due to nonlinearities in more complex PDEs. Additional details on hyperparameters and the implementation of data augmentations can be found in Appendix D.4.
@@ 3.2.1 Computer Vision Strategies
Firstly, we consider an early work that pretrains a model to verify if a video is in the correct temporal order (Misra et al., 2016). This problem is formulated as a binary classification task in which a shuffled video and the original video are assigned separate labels; within this work, we refer to this as Binary pretraining.
Subsequent work proposed methods that not only verify temporal order, but can also sort temporally shuffled video frames (Lee et al., 2017). This is generally formulated as a $n-$way classification task, where $n$ denotes the number of permutations in which a sequence of frames can be sorted. In the context of physics data, we can opt to shuffle the data spatially or temporally, as such we refer to these two pretraining strategies as TimeSort or SpaceSort. Empirically, SpaceSort does not perform well, so we omit this strategy from our results.
@@ 3.2.2 PDE Predictive Strategies
Lastly, numerical solutions of PDEs tend to leverage information of local relationships to solve equations. For example, finite difference schemes use information from neighboring nodes to calculate spatial derivatives. Motivated by this, we propose a pretraining strategy that randomly masks data in space and time and uses this incomplete information to reconstruct the full solution. This is implemented by patching the solution in space and time, randomly replacing masked patches with a learnable mask token, and regressing the true solution; we refer to this strategy as Masked.
@@ 3.2.3 Contrastive Strategies
A common strategy for pretraining in computer vision domains is to exploit similarities in the data to align samples in latent space. A proposed strategy to do this for PDEs is Physics Informed Contrastive Learning (PICL), which uses a Generalized Contrastive Loss (Leyva-Vallina et al., 2023) to cluster PDE data based on their coefficients in latent space (Lorsung & Farimani, 2024). Another strategy for self-supervised learning of PDE dynamics is using an encoder to align Lie augmented or physically invariant latent PDE samples (Mialon et al., 2023; Zhang et al., 2023b). Both works require the use of a specific encoder along with the neural operator backbone; to adapt these strategies to our experimental setup we consider directly pretraining the neural operator contrastively with these strategies. However, these methods did not seem to show significant improvements over no pretraining, as such, the results are omitted from the paper.
@@ 4.2 Neural Operators
To compare different pretraining and data augmentation strategies, we consider their effects on improving the PDE prediction performance of different neural operators. Specifically, we consider the neural operators: Fourier Neural Operator (FNO) (Li et al., 2021), DeepONet (Lu et al., 2021a) and OFormer (Li et al., 2023a). Additionally, we consider the Unet model; while it is not explicitly a neural operator, it is commonly used in literature and has shown good performance (Ronneberger et al., 2015; Gupta & Brandstetter, 2022). These neural operators are first trained using a pretraining strategy before being fine-tuned on a PDE prediction task; this could either be fixed-future prediction to model a static solution or autoregressive prediction to model a time-dependent solution. In all experiments, prediction tasks are formulated using only solution field values and grid information. Additional details on the model hyperparameters and implementation can be found in Appendix D.2.
@@ 5.2 Comparison of Data Augmentations
To study the effects of augmenting data during pretraining and finetuning, we conduct experiments in which data augmentations are added to three pretraining strategies (None, Transfer, PICL). These experiments are run to compare data augmentations to a baseline model that is not pretrained, as well as its effects on the most effective pretraining strategies (i.e. Transfer, PICL). The results are summarized in Table 3(b) for auto-regressive prediction, and the complete results can be found in Appendix B. We find the best augmentation by considering the pretraining strategy and augmentation pairing with the lowest error. To calculate its improvement, this error is compared to a model that is not pretrained.
@@
(a) Pretraining is more beneficial when downstream data is scarce. Improvement is measured by comparing the best pretraining method with no pretraining. We report the average improvement across the Heat, Adv, and Burgers PDEs for a given # of samples and model.
@@
(b) Pretraining is more beneficial when downstream data is similar to pretraining data. For a given distribution of downstream data, we compare the best pretraining strategy with no pretraining and average across PDEs in the distribution. The models show different generalization capacities.
@@
We find that different models benefit from different augmentations; for example, DeepONet performs well with shifted data, but OFormer performs well with noised data. However, across models, datasets, and downstream tasks, one can generally find a data augmentation that improves performance. This suggests that the most effective pretraining frameworks should incorporate a data augmentation strategy, and indeed the best-performing models considered in this study often make use of data augmentations. Transfer learning performs best in nine of our 12 cases, and shift augmentation performs best in eight of our 12 cases, with their combination performing best in six, suggesting that this combination improves performance best across different data sets and models. We believe that data augmentations can help due to the fact that PDE data remains scarce; numerical simulation is needed for high quality data, and as a result emulating a larger dataset with augmentations is beneficial.

### v2

@@ Abstract
Pretraining for partial differential equation (PDE) modeling has recently shown promise in scaling neural operators across datasets to improve generalizability and performance. Despite these advances, our understanding of how pretraining affects neural operators is still limited; studies generally propose tailored architectures and datasets that make it challenging to compare or examine different pretraining frameworks. To address this, we compare various pretraining methods without optimizing architecture choices to characterize pretraining dynamics on different models and datasets as well as to understand its scaling and generalization behavior. We find that pretraining is highly dependent on model and dataset choices, but in general transfer learning or physics-based pretraining strategies work best. In addition, pretraining performance can be further improved by using data augmentations. Lastly, pretraining can be additionally beneficial when fine-tuning in scarce data regimes or when generalizing to downstream data similar to the pretraining distribution. Through providing insights into pretraining neural operators for physics prediction, we hope to motivate future work in developing and evaluating pretraining methods for PDEs.
11footnotetext: Equal Contribution. Corresponding Author.
@@ 1 Introduction
Neural operators for PDEs have gained substantial interest in recent years due to their ability to quickly predict physics through inference (Li et al., 2021; Lu et al., 2021a; Brandstetter et al., 2023). Despite potential speed gains, neural operators currently struggle to generalize to unseen physics, and initial training can be slow (Lu et al., 2022; Gupta & Brandstetter, 2022). To address this issue, many works have explored different strategies to improve generalization by incorporating additional system information (Lorsung et al., 2024; Liu et al., 2023b; Takamoto et al., 2023) and pretraining neural operators across large, diverse physics to quickly fine-tune to solve PDEs (Hao et al., 2024; McCabe et al., 2023; Shen et al., 2024; Hang et al., 2024; Goswami et al., 2022). Despite showing good performance, these works usually require the use of tailored neural operators and datasets to learn different physics. This contrasts with broader deep learning trends in which pretraining methods can universally benefit models; for example, pretraining losses that are applied across CNN models (Noroozi & Favaro, 2016; Lee et al., 2017) or GNN models (Hu et al., 2020). As a result, in this work, we consider existing pretraining frameworks, as well as propose novel methods for pretraining PDE models that are flexible and can be applied across architectures or datasets.
@@
Through this work, we hope to broaden the understanding of how neural operators can be pretrained for physics prediction. We organize existing pretraining strategies, propose novel vision-inspired strategies, and include common pretraining baselines to assemble a broad set of methods for learning PDE representations. We focus on vision-inspired strategies because the point-wise representation of system values on a grid for a PDE is analogous to pixel-based representations of images, and the temporal structure of PDE system evolution is analogous to videos. We find that PDE pretraining varies depending on model and dataset choice, but in general using transfer or physics-based pretraining strategies work well. In addition, transformer or CNN-based architectures tend to benefit more from pretraining than vanilla neural operators. Furthermore, the use of data augmentations consistently improves pretraining performance in different models, datasets, and pretraining strategies. Lastly, we find that pretraining is more beneficial when fine-tuning in low-data regimes, or when downstream data is more similar to pretraining data. We hope that these insights can be used to guide future work in the development and evaluation of pretraining methods for PDEs. We make code available here: https://github.com/anthonyzhou-1/pretraining_pdes , and use 2D PDE datasets from Zhou & Farimani (2024b) which can be found here: https://zenodo.org/records/13355846 .
@@ 2 Related Works
The field of neural operators has grown rapidly in recent years, with many architectures developed to accurately approximate solutions to PDEs (Li et al., 2021; 2023a; Gupta & Brandstetter, 2022; Brandstetter et al., 2023; Lu et al., 2021a). Many works expanded on this to propose architectures to solve PDEs more quickly, with less compute, or on irregular grids (Li et al., 2023b; Hemmasian & Barati Farimani, 2023; Li et al., 2023c), and as a result, within a range of test problems, neural operators can solve PDEs quickly and accurately. However, neural operators still struggle to generalize across diverse physics, and as a result many approaches have been developed to pretrain neural operators. We summarize these past works in Table 1, and briefly describe the main approaches here. Furthermore, we review additional work from the broader field of pretraining and transfer learning due to its relevance to PDE problems.
Category PDEs Characteristic Reference
Transfer Darcy, Elasticity Fine-tuning task layers to transfer between domains Goswami et al. (2022)
Poisson, INS, Wave, FP Desiging a transferrable model by modifying neurons Zhang et al. (2023c)
Heat, Adv, Nag, Burg, NS, AC Combining operator modules with gating Tripura & Chakraborty (2023)
NS, Elasticity Pos. encoding and masked pretraining across PDEs Rahman et al. (2024)
Large Models Poisson, Helm Scaling model and datasets to characterize transfer Subramanian et al. (2023)
CNS, SWE, DiffReact Embedding PDEs to a common space, with Axial ViT McCabe et al. (2023)
INS, CNS, SWE, DiffReact Denoising & Fourier attention with large models/data Hao et al. (2024)
INS, CNS, SWE, DiffReact Conditional transformer across large PDE datasets Hang et al. (2024)
Poisson, Helm, NS, Wave, AC Scaling operator transformers to large datasets Herde et al. (2024)
Contrastive KdV, Burg, KS, INS Lie Symmetries for contrastive learning Mialon et al. (2023)
Heat, Advection, Burg Physics-informed distance in a contrastive framework Lorsung & Farimani (2024)
Burg, Adv-Diff, NS Physical invariances to contrastively learn an encoder Zhang et al. (2023b)
Meta-Learning HGO, Elasticity, Tissue Model-agnostic meta-learning loss across tasks Zhang et al. (2023a)
LV, GS, NS Novel loss term to maximize learning between PDEs Yin et al. (2021)
LV, GS, GO, NS Hyper-network to adapt operators for specific tasks Kirchmeyer et al. (2022)
In-Context Poisson, Helm, DiffReact, NS Masked pretraining and in-context learning for PDEs Chen et al. (2024)
Poisson, DiffReact In-context learning by prompting a transformer Yang et al. (2023)
@@ 2.2 Large PDE Modeling
An extension of transfer learning is to train large models on diverse physics datasets, with the intention of learning transferable representations through scaling behavior (Wei et al., 2022; Kaplan et al., 2020; Brown et al., 2020). 61 initially explores this scaling behavior by training large neural operator models on large PDE datasets to evaluate its ability to adapt to different coefficients. McCabe et al. (2023) propose a tailored architecture for solving problems across different physics, and Hao et al. (2024) expand on this by making architectural advancements and training on more diverse physics. Despite different approaches and datasets, these works generally rely on tailored, scalable architectures for large PDE datasets; pretraining is framed as physics prediction across diverse physics and fine-tuning is done on the pretraining distribution or on unseen coefficients/PDEs.
@@ 2.4 Meta/In-context Learning for PDEs
2.5 Pretraining and Transfer Learning
While PDE problems contain unique structure and challenges, works from the broader field of pretraining and transfer learning can still be leveraged to inform hypotheses and inspire novel work. In particular, many previous works seek to understand how models can learn representations that transfer across tasks or domains, as well as design architectures and pretraining tasks to optimize this (Jiang et al., 2022; Liu et al., 2023a; Zhuang et al., 2020). Under this framework, most previous work in the PDE domain is considered supervised pretraining, with the notable exception of contrastive methods. Additionally, novel works tend to focus on improving PDE-specific architectures, rather than designing specific objectives for task or domain transfer. We hypothesize that this is a reflection of the need for numerical accuracy in physics problems; prior work generally focuses on empirically improving prediction accuracy on different benchmark PDEs rather than leveraging theoretical justifications. This could highlight potential future opportunities to examine more complex or principled pretraining learning methods such as adversarial learning or causal learning.
@@ 3.1.2 Physics-Agnostic Data Augmentations
In computer vision literature, many successful data augmentations heavily modify inputs (Chen et al., 2020); in particular, cropping and cutting out portions of an image would not respect physics if adapted to the PDE domain. Following this, we investigate the effect of data augmentations that are physics-agnostic, in that they can be applied to any PDE since the augmentation does not preserve the underlying dynamics. Following recent work on denoising neural operator architectures (Hao et al., 2024), we consider adding Gaussian noise during pretraining (Noise). Furthermore, we consider scaling the PDE solution (Scale), in which the PDE solution values are scaled by a random constant, an approach similar to a color distortion, whereby the hue, contrast, saturation, or lightness of an image is scaled. Although these visual characteristics do not have direct counterparts in the PDE domain, their scaling can interpreted as scaling the intensity of a PDE solution. For certain simple PDEs, scaling can preserve physics, but this is not generally true due to nonlinearities in more complex PDEs, such as the Burgers equation or Navier-Stokes equations. Additional details on hyperparameters and the implementation of data augmentations can be found in Appendix D.4.
@@ 3.2.1 Computer Vision Strategies
Firstly, we consider an early work that pretrains a model to verify if a video is in the correct temporal order (Misra et al., 2016). This problem is formulated as a binary classification task in which a shuffled video and the original video are assigned separate labels; within this work, we refer to this as Binary pretraining. Subsequent work proposed methods that not only verify temporal order, but can also sort temporally shuffled video frames (Lee et al., 2017). This is generally formulated as a $n-$way classification task, where $n$ denotes the number of permutations in which a sequence of frames can be sorted. In the context of physics data, we can opt to shuffle the data spatially or temporally, as such we refer to these two pretraining strategies as TimeSort or SpaceSort. Empirically, SpaceSort does not perform well, so we omit this strategy from our results and analysis. For completeness, we include a preliminary set of results demonstrating this in Appendix E.
@@ 3.2.2 PDE Predictive Strategies
Lastly, numerical solutions of PDEs tend to leverage information of local relationships to solve equations. For example, finite difference schemes use information from neighboring nodes to calculate spatial derivatives. Motivated by this, we propose a pretraining strategy that randomly masks data in space and time and uses this incomplete information to reconstruct the full solution. This is implemented by patching the solution in space and time, randomly replacing masked patches with a learnable mask token, and regressing the true solution; we refer to this strategy as Masked. We adapt mask tokens to be learnable in physics space, rather than embedding space, in order to be used by any model beyond just transformers.
@@ 3.2.3 Contrastive Strategies
A common strategy for pretraining in computer vision domains is to exploit similarities in the data to align samples in latent space. A proposed strategy to do this for PDEs is Physics Informed Contrastive Learning (PICL), which uses a Generalized Contrastive Loss (Leyva-Vallina et al., 2023) to cluster PDE data based on their coefficients in latent space (Lorsung & Farimani, 2024). Another strategy for self-supervised learning of PDE dynamics is using an encoder to align Lie augmented or physically invariant latent PDE samples (Mialon et al., 2023; Zhang et al., 2023b). Both works require the use of a specific encoder along with the neural operator backbone, since a separate model is contrastively pretrained to provide conditioning information to a backbone for physics prediction. To adapt these strategies to our experimental setup, we consider directly pretraining the neural operator contrastively with these strategies. However, these methods did not seem to show significant improvements over no pretraining, as such, they are omitted from our main results and analysis. Interested readers are directed to Appendix E for preliminary results regarding these methods.
@@ 4.2 Neural Operators
To compare different pretraining and data augmentation strategies, we consider their effects on improving the PDE prediction performance of different neural operators. Specifically, we consider the neural operators: Fourier Neural Operator (FNO) (Li et al., 2021), DeepONet (Lu et al., 2021a) and OFormer (Li et al., 2023a). Additionally, we consider the Unet model; while it is not explicitly a neural operator, it is commonly used in literature and has shown good performance (Ronneberger et al., 2015; Gupta & Brandstetter, 2022). These neural operators are first trained using a pretraining strategy before being fine-tuned on a PDE prediction task; this could either be fixed-future prediction to model a static solution or autoregressive prediction to model a time-dependent solution. In all experiments, prediction tasks are formulated using only solution field values and grid information. Additional details on the model hyperparameters and implementation can be found in Appendix 5.
@@ 5.2 Comparison of Data Augmentations
To study the effects of augmenting data during pretraining and finetuning, we conduct experiments in which data augmentations are added to three pretraining strategies (None, Transfer, PICL). These experiments are run to compare data augmentations to a baseline model that is not pretrained, as well as its effects on the most effective pretraining strategies (i.e. Transfer, PICL). The results are summarized in Table 3(b) for auto-regressive prediction, and the complete results can be found in Appendix B. We find the best augmentation by considering the pretraining strategy and augmentation pairing with the lowest error. To calculate its improvement, this error is compared to a model that is not pretrained, however, to isolate the effect of pretraining, this can be compared to a corresponding pretrained model without data augmentation in the Appendix.
We find that different models benefit from different augmentations; for example, DeepONet performs well with shifted data, but OFormer performs well with noised data. However, across models, datasets, and downstream tasks, one can generally find a data augmentation that improves performance. This suggests that the most effective pretraining frameworks should incorporate a data augmentation strategy, and indeed the best-performing models considered in this study often make use of data augmentations. Furthermore, when an augmented model is compared to its baseline, the performance tends to increase even when the baseline is pretrained, albeit modestly. Transfer learning performs best in nine of our 12 cases, and shift augmentation performs best in eight of our 12 cases, with their combination performing best in six, suggesting that this combination improves performance best across different data sets and models. We believe that data augmentations can help due to the fact that PDE data remains scarce; numerical simulation is needed for high quality data, and as a result emulating a larger dataset with augmentations is beneficial.
@@
(a) Pretraining is more beneficial when downstream data is scarce. Improvement is measured by comparing the best pretraining method with no pretraining. We report the average improvement across the Heat, Adv, and Burgers PDEs for a given # samples and model.
@@
(b) Pretraining improves performance for both in and out-of-distribution downstream tasks. For each distribution of downstream data, we find the best improvement from pretraining and average across PDEs. Models show varying generalization capacities.